OpenStudy (anonymous):

x+3y=7 2x+2y=6 solve... please help

OpenStudy (anonymous):

x=-1 y=-2

OpenStudy (anonymous):

x+3y=7 ---->A 2x+2y=6---->B from A we have x = 7-3y put this in B 2(7-3y) +2y =6 solve this now

OpenStudy (anonymous):

\[x+3y=7[1]\] \[2x+2y=6[2]\] \[[1]\] \[x=7-3y\] \[[2]\] \[2(7-3y)+2y=6\] \[14-6y+2y=6\] \[14-4y=6\] Make y the subject and solve.